A film distribution manager calculates that 9% of the films released are flops. If the manager is correct, what is the probability that the proportion of flops in a sample of 407 released films would differ from the population proportion by more than 3%? Round your answer to four decimal places.

Answer: 0.0348
Answer #1

The population mean of the proportion is Mp = 0.09 .

The standard deviation of the the proportion is Op = Vn p(1-p)_ = 0.09(1 – 0.09) - 0.0142 407

The sample proportion is approximately normally distributed. P~ N(0.09,0.01423) .

The required probability,

0.03 0.03 plo<0.09 -0.03UP >0.09+ 0.03) = P(Z <-ot -0.014220.0142) 1 -26 0.03 P <0.09 -0.03UP >0.09+0.03) = 0.0142 1 0.03 P <